OpenStudy (alina123):

In a comparison taste test of two juice drinks, 780 people preferred Cranberry Blast. Only 220 people preferred Melon Splash. Complete each statement. a. There were ______ more people who preferred Cranberry Blast. b. In the taste test, ______ % of the people preferred Cranberry Blast. c. People who preferred Cranberry Blast outnumbered those who preferred Melon Splash by a ration of _ to _ Please Tell me how you got your answer by showing work or explaining.

OpenStudy (anonymous):

What's lucky for you is that 1000 people were tested. That makes things really easy 780 + 220 = 1000

OpenStudy (anonymous):

If 780 out of 1000 preferred Cranberry, then 780/1000 = 78/100 = 78% Is that clear?

OpenStudy (alina123):

SO a is 78. and b is 78%?

OpenStudy (anonymous):

No that is B. B is a percent problem A asks you how many more is 780 than 220. 780 - 220 = 560

OpenStudy (anonymous):

C asks you to put the problem into a proportion. 220 / 780. Does that make sense?

OpenStudy (alina123):

How do you though that 560 people preferred more Cranberry Blast?

OpenStudy (anonymous):

Because 220 like Melon and 780 liked Cranberry. The number of people that preferred it more is the additional people between 220 and 780 To find that, I can subtract 220 from 780 (780 - 220) to get the number.

OpenStudy (alina123):

oh, but if you wanted to find out about melon then what would you do

OpenStudy (anonymous):

Well less people like Melon, so there wouldn't be a problem like that. Technically -560 people like Melon more because 220 - 780 = -560 But that is too confusing. I suggest focusing on the problem.

OpenStudy (alina123):

ok, so a is 560, b is 78%

OpenStudy (alina123):

but i have a question about c

OpenStudy (alina123):

see it says cranberry first but in a fraction you wrote melon first so i could switch the two numbers around

OpenStudy (anonymous):

And you know what, you are right. The answer is looking for a hole number, not a decimal that is less than 1. You are correct in that it would be 780/220 Or to simplify it: 39/11

OpenStudy (alina123):

thank you

OpenStudy (anonymous):

Welcome :)